# Quantum Architecture Search with Unsupervised Representation Learning

### v5

@@ A.2 Application Settings
In the field of QC, VQE [13, 12] is a hybrid quantum algorithm for quantum chemistry, quantum simulations, and optimization problems. It is used to compute the ground state energy of a Hamiltonian based on the variational principle. For the 4- and 12-qubit quantum chemistry experiment, we use the Hamiltonian of the molecule and and its common approximate ground state energy and as the optimal energy. As for the 8-qubit experiment, we consider transverse field Ising model (TFIM) with the Hamiltonian as follows:
@@
A.3 Hyperparameters of Pre-training
@@
| epochs | traning epochs | 16 |
@@
A.4 Best Candidate circuits

### v6

@@ A.2 Application Settings
In the field of quantum chemistry, VQE [13, 12] is a hybrid quantum algorithm for quantum chemistry, quantum simulations, and optimization problems. It is used to compute the ground state energy of a Hamiltonian based on the variational principle. For the 4- and 12-qubit quantum chemistry experiment, we use the Hamiltonian of the molecule and and its common approximate ground state energy and as the optimal energy. As for the 8-qubit experiment, we consider transverse field Ising model (TFIM) with the Hamiltonian as follows:
@@
A.3 Ablation Studies
In this subsection, we analyze the sensitivity of our method to key hyperparameters of the unsupervised representation learning stage, namely the latent dimension and the depth of the GIN encoder. We evaluate both representation quality and downstream QAS performance using REINFORCE-based and Bayesian Optimization–based search on the 4-qubit MaxCut task. In Table 6, we observe that the reconstruction objective (VAE loss) varies with the latent dimension and GIN depth, whereas the downstream QAS performance is much less sensitive: both REINFORCE and Bayesian Optimization achieve comparable rewards across these configurations.
| VAE loss | avg. reward (RL) | avg. reward (BO) | ||
| 8 | 5 | 0.11 | ||
| 16 | 5 | 0.09 | ||
| 32 | 5 | 0.14 | ||
| 16 | 3 | 0.10 | ||
| 16 | 7 | 0.09 |
A.4 Hyperparameters of Pre-training
@@
| epochs | training epochs | 16 |
@@
A.5 Best Candidate circuits